A team with a cost of $160,000, an estimated residual value of $ 40,000 and an estimated life of 15 years, is amortized by the straight line method for 4 years. Due to obsolescence, it was determined to shorten the remaining useful life by 3 years and change the residual value to zero. The depreciation expense for the current and future year is: $11,000 $8,000 $11,636 O $16,000
Answer #1
As multiple questions have been asked, only first 4 can be answered at a time.
1
Accumulated depreciation till year 4=(160000-40000)/15*4= 32000
Book value at the end of year 4 = 160000-32000 = 128000
Depreciation expense for current and future year=128000/8=$16000
Option 4 is correct
2
The two methods for uncollectible accounts are reserve method and direct cancellation method
Option 2 is correct
3
The formula to determine cost to depreciate is Initial cost-Residual value
Option 2 is correct
4
Double declining balance does not consider residual value
Option 3 is correct